OpenStudy (anonymous):

4. Amir pitched a baseball at an initial height of 5.9 feet with a velocity of 73.16 feet per second. About how long will it take the ball Amir pitched to hit the ground? (1 point)4.66 secs 2.94 secs 2.35 secs 1.47 secs 5. A golf ball is hit from the ground with an initial velocity of 208 feet per second. Assume the starting height of the ball is 0 feet. How long will it take the golf ball to hit the ground? (1 point)13 secs 21 secs 42 secs 6 secs 6. A particle is moving along a projectile path where the initial height is 160 feet with an initial speed of 144 feet per secon

OpenStudy (anonymous):

1.d 2.a 6. A particle is moving along a projectile path where the initial height is 160 feet with an initial speed of 144 feet per second. What is the maximum height of the particle? (1 point)128 feet 224 feet 272 feet 484 feet

OpenStudy (anonymous):

I need 4 and 5 checked over and I need help with 6.

OpenStudy (anonymous):

224 =P

OpenStudy (anonymous):

can you explain why?

OpenStudy (anonymous):

no i.....did it in my head...

OpenStudy (anonymous):

i need help with one more is that fine?

OpenStudy (anonymous):

7. Use complete sentences to explain how you would use graphing technology to find the solutions of 0 = −x2 + 2.5x + 1. What do the solutions of 0 = −x2 + 2.5x + 1 represent on the graph of the equation?

OpenStudy (anonymous):

im sorry graphing is what im worst at =[

OpenStudy (anonymous):

oh ok animalain?

OpenStudy (anonymous):

how about you?

OpenStudy (anonymous):

anyone number 7?

OpenStudy (anonymous):

I don't see how this works, exactly. For example, a pitched baseball goes approximately horizontal. I don't see any way a ball thrown that way can stay aloft for anything close to the times you have on the answers listed. On the other problems, there is speed information, but no directional information, which is probably needed to find the answers.

OpenStudy (anonymous):

Number seven I can handle..... Take the function, and put it into the computer or calculator to graph. Then look for the x values that the curve (parabola in this case) intercepts the x axis. Those numbers are the solutions of your quadratic equation.

OpenStudy (anonymous):

and that means?

OpenStudy (anonymous):

1) Tell the calculator to draw the graph of y=x^2+2.5x+1 2) Find where the graph of the function intersects the x axis. 3) Those x values are the answers.

OpenStudy (anonymous):

i will still need to explain though.

OpenStudy (anonymous):

Hello? Finally someone could help with number 7.

OpenStudy (anonymous):

I don't understand what more you need to know about number seven.

OpenStudy (anonymous):

oh i get it what you wrote above was the answer? im so stupid.

OpenStudy (anonymous):

Don't say that. Get it now?

OpenStudy (anonymous):

yes. thanks for explaining it.lol

OpenStudy (anonymous):

No sweat.